# COBASE: A new copula-based shuffling method for ensemble weather forecast postprocessing

### v1

@@ Abstract
Statistical post-processing of weather forecasts , Two-step parametric correction , Dependence modeling and copulas , ALADIN-LAEF ensemble system , ECMWF.
@@ 1 Introduction
The forecasting skills of the physical numerical weather prediction (NWP) models have been consistently improved in recent years (e.g., Bauer et al., 2015). Efforts in research have focused on enhancing physical parameterizations, assimilating non-conventional data, and boosting spatial resolution. Despite the progress made, deterministic and probabilistic NWP forecasts still show systematic errors and underdispersion. Potential causes of errors are imperfect measurements (which also serve as initial conditions, e.g., Cardinali, 2009), imprecise parametrization of physical equations (e.g., Bechtold et al., 2014) and inherent unpredictability due to chaotic dynamics in atmospheric flows (e.g., Lorenz, 1963). To account for such model errors, calibration approaches based on statistical or machine learning post-processing techniques have been developed in the last years: see, for example, Gneiting et al. (2005); Raftery et al. (2005); Möller et al. (2013); Schefzik et al. (2013); Taillardat et al. (2016); Schefzik (2016); Scheuerer and Hamill (2018); Rasp and Lerch (2018); Lerch et al. (2020); Vannitsem et al. (2021). The standard statistical methods for the post-processing of raw ensemble forecasts are Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005) and Bayesian Model Averaging (BMA) (Raftery et al., 2005). These techniques work well in practice, but are designed to improve predictions for only one variable at a time, e.g., for a fixed weather variable at a single location and specific lead-time. As a result, such methods fail to maintain the inter-variable and spatio-temporal dependencies of the forecasts (e.g., Schefzik, 2016). This poses challenges for forecasting high-impact weather events such as floods (Schaake et al., 2010), heat waves (e.g., Baran et al., 2020), or the onset of hurricanes, where numerous weather factors intricately interact. To address this problem, multivariate postprocessing methods have been developed. The most popular techniques are based on empirical copulas (e.g., Schefzik et al., 2013; Lakatos et al., 2023). Namely, these are two-step approaches that use standard EMOS or BMA to correct the univariate forecast and a reshuffling technique to reconstruct the lost dependencies in the samples that are drawn from the univariate distributions. There are several ways to sample from the post-processed distribution, including drawing random samples, uniform quantiles, or stratified samples. Uniform quantiles or stratified samples generally perform better than random samples, particularly for extremes (Wilks, 2015; Hu et al., 2016). Examples of reshuffling methods are Schaake Shuffle (Clark et al., 2004), Ensemble Copula Coupling (ECC) (Schefzik et al., 2013), and Sim Schaake (Schefzik, 2016). The common ground of these methods is that the dependence structure can be obtained from a reference sample, which can be based on past observations (for Schaake Shuffle and Sim Schaake) or on the raw forecasts (for ECC). However, these methods have some important limitations as they require a large number of past observations or a reliable, exchangeable ensemble system that simulates a realistic dependence structure to be effective. These conditions are not always satisfied as is the case for the ALADIN-LAEF system used in Austria (Wang et al., 2011). In such cases, multivariate parametric approaches that do not require extensive data archives nor do not assume exchangeability, could be better suited for multivariate post-processing.
@@
Multivariate parametric approaches for the statistical postprocessing of weather forecasts already exist (e.g., Möller et al., 2013; Schefzik and Möller, 2018; Jobst et al., 2024; Lerch et al., 2020). Parametric methods with advanced copula models improve the dependence structure of the corrected sample but often fail to maintain a suitable marginal calibration as they can only be used with random sampling which might perform poorly compared to uniform quantile sampling (Wilks, 2015). This is a problem that puts copula-based methods at a disadvantage when compared to non-parametric two-step approaches which can be applied to any univariate corrected sample. We believe that this limitation has hindered the potential of multivariate parametric postprocessing in both academic research and operational use.
@@
This paper introduces a novel approach aimed at fully harnessing the potential of two-step parametric correction techniques. Specifically, we introduce a new copula-based shuffling technique (COBASE) that leverages the adaptability of parametric copulas while preserving the marginal calibration. We present the methodology through two case studies, similar to those examined in Perrone et al. (2020) and Whan et al. (2021). Our attention is directed towards spatial dependencies in 2-meter temperature predictions using the ALADIN-LAEF system in Austria and inter-variable and spatial dependencies in 2-meter temperature and dew-point temperature predictions the Netherlands. The case studies demonstrate the applicability of our method in different climates, with different numerical weather prediction systems (regional and global) and number of dimensions. Our approach significantly outperforms the state-of-the-art multivariate parametric methods, and it provides a viable way to enhance the performance of any parametric copula-based method in the post-processing context.
@@
The paper is organized as follows. In Section 2, we present the two numerical weather prediction systems and our case study settings. In Section 3, we review existing non-parametric and parametric methods for post-processing, discuss the limitations of current approaches, and introduce COBASE, our new methodology. Section 4 tests COBASE in our case studies. Finally, we draw conclusions and end with a general discussion in Section 5.
@@ 2.2 Case study 2: inter-variable and spatial dependencies in the Netherlands
For this case study, we take operational forecasts of 2-meter temperature (T2m) and dew-point temperature (DPT) from the global ensemble prediction system run by the European Center for Medium-Range Weather Forecasting (ECMWF). We use the 51-members that are initialized each day at 12 UTC during the period 2021-2024, and take the forecasts that are valid at +24 hours. There were several model upgrades during this time, including a change in horizontal resolution from  18km to  9km in 2023. The use of a rolling training window, as explained in Section 3, reduces the impact of these model changes.
@@ 3 Methods
This section outlines the two-step approaches currently considered state-of-the-art in multivariate statistical post-processing. The key idea of these methods is to separate the correction of forecasts into two parts, one that addresses the marginal forecasts, and another that pertains the dependence structure. In practical terms, this involves initially applying univariate post-processing techniques to each margin separately, meaning that we address each weather variable, lead time, and station on an individual basis. Afterwards, we restore the missing dependencies by applying the theory of copulas.
Copulas are cumulative joint distributions with uniform margins on $[0,1]$. The theoretical foundation of copula theory is Sklar’s theorem (Sklar, 1959), which states that the $d$-dimensional multivariate cumulative distribution function $F$ of a random vector $(X_{1},\ldots,X_{d})$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as the composition of a copula $C$ and univariate marginal distributions, i.e., $F(x_{1},\ldots,x_{d})=C(F_{1}(x_{1}),\ldots,F_{d}(x_{d}))$. Therefore, copulas can identify the dependence structure of a phenomenon without considering its marginal distributions. In the post-processing context, they can be used in the second step to reintroduce a dependence structure which gets lost in the univariate post-processing and sampling step. In the remainder of the section, we first introduce the univariate correction method used in our study, followed by the nonparametric and parametric copula-based approaches, and our newly proposed method COBASE. A summary of the specifications of the methods used with the corresponding labels is given in Table 2.
3.1 Univariate Post-processing: Ensemble Model Output Statistics (EMOS)
@@
While EMOS provides effective univariate corrections, it does not capture dependencies across multiple variables or locations, motivating the use of multivariate post-processing methods. In the following section, we review non-parametric empirical copula-based postprocessing approaches.
3.2 Nonparametric multivariate post-processing: Empirical-copula methods
This section reviews the most common two-step nonparametric techniques for statistical post-processing. All of these methods are based on empirical copulas, which offer an effective way to incorporate the dependence structure into weather forecasting problems without imposing any parametric assumptions (Schefzik, 2015). An empirical copula is essentially a discrete form of a copula $C$, known as a discrete copula, derived from a given multivariate sample (Schefzik, 2015). As such, it encodes the dependence structure of the reference sample through the corresponding ranks. In weather forecasting, empirical copulas can be used to generate adjusted multivariate forecasts that integrate dependencies taken from a reference rank structure. This idea was first introduced in the Schaake Shuffle method (Clark et al., 2004), which uses the rank structure obtained from historical observations. Then, this concept was refined and new methods were developed, the most popular being ECC (Schefzik et al., 2013) and Sim Schaake (Schefzik, 2016). ECC reconstructs the dependence according to the raw ensemble forecasts. Sim Schaake combines the ideas of ECC and Schaake shuffle: First, it selects past days where the ensemble forecasts are similar to the verification day and then uses the selected observations to reconstruct the dependence.
@@ 3.2 Nonparametric multivariate post-processing: Empirical-copula methods
In addition to the construction of the reference dependence structure $DC_{ref}$, another point that can generally vary is the sampling technique used in S3 of Algorithm 1. In Schefzik et al. (2013), various sampling techniques have been discussed. One could simply take equidistant quantiles of step $N+1$ from the univariate corrected distribution $F_{l}$:
@@
S1: Derive a reference dependence structure $DC_{ref}$;
S2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
@@
S4: Shuffle the corrected univariate samples according to the ranks of $DC_{ref}$.
In our case studies, we ensure a fair comparison with ECC by setting the size of the ensemble at $N=17$ for the ALADIN-LAEF groups and $N=51$ for the ECMWF data. For the Sim Schaake and Schaake Shuffle methods, we use all available observations and raw forecasts, including past and future dates (approximately four years), to ensure a consistent number of reference dates throughout the test set. Specifically for the Schaake Shuffle, we select observations randomly from a 14-day window before and after the forecast date to preserve the seasonal climatology.
@@
    The necessity for extensive historical data to more accurately illustrate the dependence influences both the Schaake Shuffle and Sim Schaake.
@@
    The requirement for extensive historical raw forecasts is crucial for the similarity criterion, impacting Sim Schaake.
@@
    The requirement that the sample size of the corrected forecasts ($N$) must equal the number of ensemble members ($M$), imposes a significant limitation on the applicability of ECC.
Although these limitations might seem less relevant in an operational context, they may have a significant impact in practice. An example is Case Study 1 in Austria, where the post-processing of the operational system C-LAEF could not be performed due to the lack of historical raw forecasts and historical observations. Furthermore, for ALADIN-LAEF, the limited ensemble size of only $17$ members posed a substantial constraint on the application of the ECC method. Consequently, this case study demonstrates a scenario in which the shortcomings of these methods might justify a transition to alternative approaches, such as parametric multivariate postprocessing approaches. We introduce these approaches in the following section after providing an overview of how multivariate forecasts are evaluated in this study.
@@
$\text{ES}=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{m}-y\|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{\nu=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{\nu}-x_{m}\|$ (7)
@@
$VS=\sumop\displaylimits_{i=1}^{d}\sumop\displaylimits_{j=1}^{d}{w}_{i,j}\left(\left|{y}_{i}-{y}_{j}\right|^{p}-\frac{1}{m}\sumop\displaylimits_{k=1}^{m}\left|x^{(k)}_{i}-x^{(k)}_{j}\right|^{p}\right)^{2}.$ (8)
where $w_{i,j}$ are weights and $p>0$ is the order of the score. Because the VS compares pairwise differences between variables in the observation and forecast, it is particularly sensitive to errors in the dependence structure. In our case study, we use both the Energy Score and the Variogram Score to assess and compare the performance of the different multivariate post-processing methods. These metrics provide complementary information, evaluating both the overall accuracy of the multivariate forecasts and their ability to represent dependence structures. We used the variant with no weights and $p=1$ for the Variogram Scores, but we obtained consistent results for other variants as well.
3.3 Existing parametric multivariate post-processing

### v2

@@ Abstract
Statistical postprocessing of weather forecasts , Two-step parametric correction , Dependence modeling and copulas , ALADIN-LAEF ensemble system , ECMWF.
@@ 1 Introduction
The forecasting skills of the physical numerical weather prediction (NWP) models have been consistently improved in recent years (e.g., Bauer et al., 2015). Efforts in research have focused on enhancing physical parameterizations, assimilating non-conventional data, and boosting spatial resolution. Despite the progress made, deterministic and probabilistic NWP forecasts still show systematic errors and underdispersion. Potential causes of errors are imperfect measurements (which also serve as initial conditions, e.g., Cardinali, 2009), imprecise parametrization of physical equations (e.g., Bechtold et al., 2014) and inherent unpredictability due to chaotic dynamics in atmospheric flows (e.g., Lorenz, 1963). To account for such model errors, calibration approaches based on statistical or machine learning postprocessing techniques have been developed in the last years: see, for example, Gneiting et al. (2005); Raftery et al. (2005); Möller et al. (2013); Schefzik et al. (2013); Taillardat et al. (2016); Schefzik (2016); Scheuerer and Hamill (2018); Rasp and Lerch (2018); Lerch et al. (2020); Vannitsem et al. (2021). The standard statistical methods for the postprocessing of raw ensemble forecasts are Ensemble Model Output Statistics (EMOS) (Gneiting et al., 2005) and Bayesian Model Averaging (BMA) (Raftery et al., 2005). These techniques work well in practice, but are designed to improve predictions for only one variable at a time, e.g., for a fixed weather variable at a single location and specific lead-time. As a result, such methods fail to maintain the inter-variable and spatio-temporal dependencies of the forecasts (e.g., Schefzik, 2016). This poses challenges for forecasting high-impact weather events such as floods (Schaake et al., 2010), heat waves (e.g., Baran et al., 2020), or the onset of hurricanes, where numerous weather factors intricately interact. To address this problem, multivariate postprocessing methods have been developed. The most popular techniques are based on empirical copulas (e.g., Schefzik et al., 2013; Lakatos et al., 2023). Namely, these are two-step approaches that use standard EMOS or BMA to correct the univariate forecast and a reshuffling technique to reconstruct the lost dependencies in the samples that are drawn from the univariate distributions. There are several ways to sample from the post-processed distribution, including drawing random samples, uniform quantiles, or stratified samples. Uniform quantiles or stratified samples generally perform better than random samples, particularly for extremes (Wilks, 2015; Hu et al., 2016). Examples of reshuffling methods are Schaake Shuffle (Clark et al., 2004), Ensemble Copula Coupling (ECC) (Schefzik et al., 2013), and Sim Schaake (Schefzik, 2016). The common ground of these methods is that the dependence structure can be obtained from a reference sample, which can be based on past observations (for Schaake Shuffle and Sim Schaake) or on the raw forecasts (for ECC). However, these methods have some important limitations as they require on sufficiently representative past observations or a reliable, exchangeable ensemble system that simulates a realistic dependence structure to be effective. These conditions are not always satisfied as is the case for the ALADIN-LAEF system used in Austria (Wang et al., 2011). In such cases, multivariate parametric approaches that do not require extensive data archives nor do not assume exchangeability, could be better suited for multivariate postprocessing.
@@
Multivariate parametric approaches for the statistical postprocessing of weather forecasts already exist (e.g., Möller et al., 2013; Schefzik and Möller, 2018; Jobst et al., 2024; Lerch et al., 2020). Parametric methods with advanced copula models improve the dependence structure of the corrected sample but often fail to maintain a suitable marginal calibration as they can only be used with random sampling which might perform poorly compared to uniform quantile sampling (Wilks, 2015). This is a problem that puts copula-based methods at a disadvantage when compared to nonparametric two-step approaches which can be applied to any univariate corrected sample. We believe that this limitation has hindered the potential of multivariate parametric postprocessing in both academic research and operational use.
@@
This paper introduces a novel approach aimed at fully harnessing the potential of two-step parametric correction techniques. Specifically, we introduce a new copula-based shuffling technique (COBASE) that leverages the adaptability of parametric copulas while preserving the marginal calibration. We present the methodology through two case studies, similar to those examined in Perrone et al. (2020) and Whan et al. (2021). Our attention is directed towards spatial dependencies in 2-meter temperature predictions using the ALADIN-LAEF system in Austria and inter-variable and spatial dependencies in 2-meter temperature and dew-point temperature predictions the Netherlands. The case studies demonstrate the applicability of our method in different climates, with different numerical weather prediction systems (regional and global) and number of dimensions. Our approach significantly outperforms the state-of-the-art multivariate parametric methods, and it provides a viable way to enhance the performance of any parametric copula-based method in the postprocessing context.
@@
The paper is organized as follows. In Section 2, we present the two numerical weather prediction systems and our case study settings. In Section 3, we review existing nonparametric and parametric methods for postprocessing, discuss the limitations of current approaches, and introduce COBASE, our new methodology. Section 4 tests COBASE in our case studies. Finally, we draw conclusions and end with a general discussion in Section 5.
@@ 2.2 Case study 2: inter-variable and spatial dependencies in the Netherlands
For this case study, we take operational forecasts of 2-meter temperature (T2m) and dew-point temperature (DPT) from the global ensemble prediction system run by the European Center for Medium-Range Weather Forecasting (ECMWF). We use the 51-member ECMWF ensemble, consisting of one unperturbed control forecast and 50 perturbed ensemble members, that are initialized each day at 12 UTC during the period 2021-2024, and take the forecasts that are valid at +24 hours. The ensemble thus comprises two exchangeable groups: the statistically equivalent perturbed members and a distinct control member. There were several model upgrades during this time, including a change in horizontal resolution from  18km to  9km in 2023. The use of a rolling training window, as explained in Section 3, reduces the impact of these model changes.
@@ 3 Methods
This section outlines the two-step approaches currently considered state-of-the-art in multivariate statistical postprocessing. The key idea of these methods is to separate the correction of forecasts into two parts, one that addresses the marginal forecasts, and another that pertains the dependence structure. In practical terms, this involves initially applying univariate postprocessing techniques to each margin separately, meaning that we address each weather variable, lead time, and station on an individual basis. Afterwards, we restore the missing dependencies by applying the theory of copulas.
Copulas are cumulative joint distributions with uniform margins on $[0,1]$. The theoretical foundation of copula theory is Sklar’s theorem (Sklar, 1959), which states that the $d$-dimensional multivariate cumulative distribution function $F$ of a random vector $(X_{1},\ldots,X_{d})$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as the composition of a copula $C$ and univariate marginal distributions, i.e., $F(x_{1},\ldots,x_{d})=C(F_{1}(x_{1}),\ldots,F_{d}(x_{d}))$. Therefore, copulas can identify the dependence structure of a phenomenon without considering its marginal distributions. In the postprocessing context, they can be used in the second step to reintroduce a dependence structure which gets lost in the univariate postprocessing and sampling step. In the remainder of the section, we first introduce the univariate correction method used in our study, followed by the nonparametric and parametric copula-based approaches, and our newly proposed method COBASE. A summary of the specifications of the methods used with the corresponding labels is given in Table 2.
3.1 Univariate postprocessing: Ensemble Model Output Statistics (EMOS)
@@
While EMOS provides effective univariate corrections, it does not capture dependencies across multiple variables or locations, motivating the use of multivariate postprocessing methods. In the following section, we review nonparametric empirical copula-based postprocessing approaches.
3.2 Nonparametric multivariate postprocessing: Empirical-copula methods
This section reviews the most common two-step nonparametric techniques for statistical postprocessing. All of these methods are based on empirical copulas, which offer an effective way to incorporate the dependence structure into weather forecasting problems without imposing any parametric assumptions (Schefzik, 2015). An empirical copula is essentially a discrete form of a copula $C$, known as a discrete copula, derived from a given multivariate sample (Schefzik, 2015). As such, it encodes the dependence structure of the reference sample through the corresponding ranks. In weather forecasting, empirical copulas can be used to generate adjusted multivariate forecasts that integrate dependencies taken from a reference rank structure. This idea was first introduced in the Schaake Shuffle method (Clark et al., 2004), which uses the rank structure obtained from historical observations. Then, this concept was refined and new methods were developed, the most popular being ECC (Schefzik et al., 2013) and Sim Schaake (Schefzik, 2016). ECC reconstructs the dependence according to the raw ensemble forecasts. Sim Schaake combines the ideas of ECC and Schaake Shuffle: First, it selects past days where the ensemble forecasts are similar to the verification day and then uses the selected observations to reconstruct the dependence.
@@ 3.2 Nonparametric multivariate postprocessing: Empirical-copula methods
In addition to the construction of the reference dependence structure $DC_{ref}$, another point that can generally vary is the sampling technique used in Step3 of Algorithm 1. In Schefzik et al. (2013), various sampling techniques have been discussed. One could simply take equidistant quantiles of step $N+1$ from the univariate corrected distribution $F_{l}$:
@@
Step1: Derive a reference dependence structure $DC_{ref}$;
Step2: Apply EMOS and obtain a corrected distribution $F_{l}$ for each margin $l=1,\ldots,d$;
@@
Step4: Shuffle the corrected univariate samples according to the ranks of $DC_{ref}$.
In our case studies, we ensure a fair comparison with ECC by setting the size of the ensemble at $N=17$ for the ALADIN-LAEF groups and $N=51$ for the ECMWF data. For the Sim Schaake and Schaake Shuffle methods (in both datasets), we use all available observations and raw forecasts, including past and future dates after the validation date (approximately four years), to ensure a consistent number of reference dates throughout the test set. Although future observations are of course not available in an operational setting, we consider it reasonable to assume that four years of past observations and forecasts can be accessed, which is why we include future dates in this analysis. Specifically for the Schaake Shuffle, we select observations randomly from a 14-day window before and after the forecast date to preserve the seasonal climatology. We additionally examined the performance using larger window sizes and observed no notable differences in the outcomes.
@@
    The direct reliance on past observations as dependence templates for reconstructing multivariate dependence, which affects both the Schaake Shuffle and Sim Schaake. This restricts the reconstructed dependence to patterns directly observed in the historical record and implicitly assumes that suitable historical analogues exist for the current forecast situation.
@@
    The additional requirement of historical raw forecasts and a sufficiently large pool of past observations/forecasts to define a meaningful similarity criterion, which specifically impacts Sim Schaake.
@@
    The requirement that the sample size of the corrected forecasts ($N$) must equal the number of ensemble members ($M$), which imposes a significant limitation on the applicability of ECC.
Although these limitations might seem less relevant in an operational context, they may have a significant impact in practice. An example is Case Study 1 in Austria, where the postprocessing of the operational system C-LAEF could not be performed due to the lack of historical raw forecasts and historical observations. Furthermore, for ALADIN-LAEF, the limited ensemble size of only $17$ members posed a substantial constraint on the application of the ECC method. Consequently, this case study demonstrates a scenario in which the shortcomings of these methods might justify a transition to alternative approaches, such as parametric multivariate postprocessing approaches. We introduce these approaches in the following section after providing an overview of how multivariate forecasts are evaluated in this study.
@@
$\text{ES}=\dfrac{1}{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{m}-y\|-\dfrac{1}{2M^{2}}\sumop\displaylimits\limits_{\nu=1}^{M}\sumop\displaylimits\limits_{m=1}^{M}\|x_{\nu}-x_{m}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}.}$ (7)
@@
$VS=\sumop\displaylimits_{i=1}^{d}\sumop\displaylimits_{j=1}^{d}{w}_{i,j}\left(\left|{y}_{i}-{y}_{j}\right|^{p}-\frac{1}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}\sumop\displaylimits_{k=1}^{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}}\left|x^{(k)}_{i}-x^{(k)}_{j}\right|^{p}\right)^{2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}$ (8)
where $w_{i,j}$ are weights and $p>0$ is the order of the score. Because the VS compares pairwise differences between variables in the observation and forecast, it is particularly sensitive to errors in the dependence structure. In our case study, we use both the Energy Score and the Variogram Score to assess and compare the performance of the different multivariate postprocessing methods. These metrics provide complementary information, evaluating both the overall accuracy of the multivariate forecasts and their ability to represent dependence structures. We used the variant with no weights and $p=1$ for the Variogram Scores, but we obtained consistent results for other variants as well.
Method Approach Dependence No direct use of past dependence templates No use of historical raw forecasts Flexible marginal sampling No restriction on the sample size
Sim Schaake Shuffling Empirical copula (from past obs) × ×
Schaake Shuffle Shuffling Empirical copula (from past obs) ×
ECC Shuffling Empirical copula (from raw forecasts) ×
Table 1: Summary of the multivariate postprocessing methods analyzed in this work and their main characteristics. Check marks indicate that a method satisfies the corresponding property, while crosses indicate that it does not.
3.3 Existing parametric multivariate postprocessing